### v2

# On a discrete version of the position-momentum commutation relation

Abstract
The usual position-momentum commutation relation plays a fundamental role in the mathematical description of continuous-variable quantum systems. In the case of a qudit described by a Hilbert space of a high enough dimension, there exists a discrete version of the position-momentum commutation relation approximately satisfied by a large part of the pure quantum states. Our purpose is to explore in more details the set of these states. We show that it contains a family of discrete-variable Gaussian states depending on a continuous parameter and certain discrete coherent states. It also contains various discrete-variable versions of the Hermite-Gauss states, defined either as eigenstates of certain discrete versions of the harmonic oscillator Hamiltonian or generated by using a discrete version of the creation or annihilation operator. As a direct consequence, a discrete version of the incertitude relation is satisfied by the considered quantum states.
preprint: APS/123-QED
I Introduction
The quantum systems with finite-dimensional Hilbert space [1, 2, 3, 4, 5] play an important role in the investigation of quantum mechanics foundations, in Quantum Information, and in other applications. The mathematical description of these discrete-variable systems is obtained by following, as much as possible, the analogy with the continuous-variable quantum systems. In certain cases, the discrete-variable versions have properties very similar to those concerning the continuous systems, but this does not happen in all the cases.
In quantum mechanics, a pair $A$ and $B$ of Hermitian operators on a Hilbert space is canonical if they satify the relation $A,B]\!=\!\mathrm{i}\hbar$, where [A,B]=ABBA[A,B]\!=\!AB-\!BA, $h\!=\!2\pi\hbar$ is the Plank constant, and $\mathrm{i}\hbar\equiv\mathrm{i}\hbar I$ is the identity operator multiplied by $\mathrm{i}\hbar$. In the finite-dimensional case, such a pair does not exist because the trace of the two sides of $A,B]\!=\!\mathrm{i}\hbar$ is different. In [6], the authors show that certain canonical pairs can be obtained by relaxing the definition, namely by imposing the relation $A,B]\!=\!\mathrm{i}\hbar$ to be satisfied only on a proper subspace of the Hilbert space. Some references concerning possible applications of this new definition are presented in [6].
Our approach is different. We show that a discrete version of the position-momentum commutation relation $\bar{q},\bar{p}]\!=\!\mathrm{i}\hbar$ is approximately satisfied on a proper subspace if the dimension of the Hilbert space is high enough. Among the elements of such a subspace there are some of the most important states used in the mathematical description of qudits [7, 8, 9, 10].
The presented results may be used as a starting point of new mathematical formalisms involving only certain particular states of a qudit, by following, for example, the analogy with the theory of continuous-variable Gaussian states used in Quantum Information [11, 12, 13].
II Quantum systems with finite-dimensional Hilbert space
A spinless particle with mass, in one dimension, is usually described by using the Hilbert space
$L^{2}(\mathbb{R})\!=\!\overline{\left\{\Psi\!:\!\mathbb{R}\!\rightarrow\!\mathbb{C}\ \left|\ \int_{-\infty}^{\infty}|\Psi(q)|^{2}\,dq\!<\!\infty\right.\right\}}$ (1)
(closure of the space of square integrable functions) with
$\left\langle\Phi,\Psi\right\rangle\!=\!\int_{-\infty}^{\infty}\overline{\Phi(q)}\,\Psi(q)\,dq.$ (2)
In order to have a simple analogy, we shall consider only the case of a discrete-variable quantum system described by an odd-dimensional (d=2s+1d\!=\!2s\!+\!1) Hilbert space $\mathcal{H}$ regarded as a space of functions, namely
$\mathcal{H}\!=\!\{\psi\!:\!\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}\!\rightarrow\!\mathbb{C}\,\}$ (3)
with the inner product
$\langle\varphi|\psi\rangle\!=\!\sum\limits_{n=-s}^{s}\overline{\varphi(n)}\,\psi(n)$ (4)
and the norm
$\psi||\!=\!\sqrt{\langle\psi|\psi\rangle}\!=\!\sqrt{\sum\limits_{n=-s}^{s}|\psi(n)|^{2}}.$ (5)
The discrete $\delta$ functions $\delta_{-s}$, $\delta_{-s+1}$, … , $\delta_{s}$, where
[…]
form an orthonormal basis of $\mathcal{H}$. By using Dirac notation $|m\rangle$ instead of $|\delta_{m}\rangle$, we have
$\langle n|m\rangle\!=\!\delta_{nm},\quad\mbox{and}\quad\sum_{m=-s}^{s}|m\rangle\langle m|\!=\!\mathbb{I},$ (7)
where $\mathbb{I}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},\ \mathbb{I}\psi\!=\!\psi$, is the identity operator.
III Discrete Fourier transform
The usual Fourier transform
$\begin{array}[]{l}\hat{F}\!:\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{F}\Psi,\\[5.69054pt] (\hat{F}\Psi)(p)\!=\!\frac{1}{\sqrt{h}}\int\limits_{-\infty}^{\infty}\mathrm{Exp}[-\frac{2\pi\mathrm{i}}{h}pq]\,\Psi(q)\,dq,\end{array}$ (8)
is a unitary transform. Its inverse is the adjoint transform
$\begin{array}[]{l}\hat{F}^{\dagger}\!:\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{F}^{\dagger}\Psi,\\[5.69054pt] (\hat{F}^{\dagger}\Psi)(p)\!=\!\frac{1}{\sqrt{h}}\int\limits_{-\infty}^{\infty}\mathrm{Exp}[\frac{2\pi\mathrm{i}}{h}pq]\,\Psi(q)\,dq,\end{array}$ (9)
The corresponding discrete-variable version [2, 7],
$\begin{array}[]{l}\hat{\mathfrak{F}}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\hat{\mathfrak{F}}\psi,\\ (\hat{\mathfrak{F}}\psi)(k)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n=-s}^{s}\mathrm{Exp}[-\frac{2\pi\mathrm{i}}{d}kn]\,\psi(n),\end{array}$ (10)
is also a unitary transform. Its inverse is the adjoint
$\begin{array}[]{l}\hat{\mathfrak{F}}^{\dagger}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\hat{\mathfrak{F}}^{\dagger}\psi,\\ (\hat{\mathfrak{F}}^{\dagger}\psi)(k)\!=\!\mbox{\small$\frac{1}{\sqrt{d}}$}\!\sum\limits_{n=-s}^{s}\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}kn]\,\psi(n).\end{array}$ (11)
IV Discrete version of position and momentum operators
The position operator
$\begin{array}[]{l}\hat{q}\!:\!D_{q}\!\subset\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{q}\,\Psi,\\[2.84526pt] (\hat{q}\,\Psi)(q)\!=\!q\,\Psi(q),\end{array}$ (12)
whose domain of definition is
[…]
admits the discrete-variable version [1, 2]
$\hat{\mathfrak{q}}:\mathcal{H}\!\rightarrow\!\mathcal{H}\!:\!\psi\mapsto\hat{\mathfrak{q}}\psi,\quad(\hat{\mathfrak{q}}\psi)(n)\!=\!n\,\psi(n).$ (14)
The basis $\{|n\rangle\}_{n=-s}^{s}$ is an eigenbasis of $\hat{\mathfrak{q}}$, and $\hat{\mathfrak{q}}$ admits the spectral resolution
$\hat{\mathfrak{q}}\!=\!\sum\limits_{n=-s}^{s}n|n\rangle\langle n|.$ (15)
The momentum operator
$\begin{array}[]{l}\hat{p}\!:\!D_{p}\!\subset\!L^{2}(\mathbb{R})\!\rightarrow\!L^{2}(\mathbb{R})\!:\!\Psi\!\mapsto\hat{p}\,\Psi,\\[2.84526pt] \hat{p}\,\Psi\!=\!-\mathrm{i}\hbar\frac{\mathrm{d}}{\mathrm{d}q}\Psi,\end{array}$ (16)
whose domain of definition is
[…]
satisfies the relation
$\hat{p}\!=\!\hat{F}^{\dagger}\hat{q}\hat{F}$. (18)
The discrete-variable version of momentum is defined as [1, 2]
$\hat{\mathfrak{p}}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},\quad\hat{\mathfrak{p}}\!=\!\hat{\mathfrak{F}}^{\dagger}\hat{\mathfrak{q}}\hat{\mathfrak{F}}$. (19)
Explicitly,
$\begin{array}[]{l}(\hat{\mathfrak{p}}\psi)(n)\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}\sum\limits_{m=-s}^{s}k\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}k(n-m)]\,\psi(m),\end{array}$ (20)
and $\hat{\mathfrak{p}}$ admits the spectral resolution
$\hat{\mathfrak{p}}\!=\!\sum\limits_{n=-s}^{s}n\,\hat{\mathfrak{F}}^{\dagger}|n\rangle\langle n|\hat{\mathfrak{F}}$, (21)
that is $\{\hat{\mathfrak{F}}^{\dagger}|n\rangle\}_{n=-s}^{s}$ is an eigenbasis of $\hat{\mathfrak{p}}$.
V Discrete version of the position-momentum commutation relation
Table 1: Eigenvalues of $\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$ in two cases.
Case $d=11$ Case $d=31$
[…]
In the continuous case, the usual position and momentum satisfy the remarkable relation
$\hat{q},\hat{p}]={\rm i}\mbox{\small$\frac{h}{2\pi}$}.$ (22)
This means that
$\hat{q}(\hat{p}\Psi)-\hat{p}(\hat{q}\Psi)\!=\!{\rm i}\mbox{\small$\frac{h}{2\pi}$}\Psi$ (23)
for any $\Psi$ from the subspace
$\left\{\,\Psi\!\in\!D_{q}\cap D_{p}\ \ |\ \ \hat{q}\Psi\!\in\!D_{p}\ \ \mbox{and}\ \ \hat{p}\Psi\!\in\!D_{q}\,\right\}.$ (24)
Table 2: Case d=11d\!=\!11: $\varphi_{1},\,\varphi_{2},...,\varphi_{d}$ are eigenstates of $\mathfrak{F}$.
$k$ $\parallel\mathfrak{F}\varphi_{k}\!-\!\varphi_{k}\parallel$ $\parallel\mathfrak{F}\varphi_{k}\!+\!\varphi_{k}\parallel$ $\parallel\mathfrak{F}\varphi_{k}\!-\!\mathrm{i}\varphi_{k}\parallel$ $\parallel\mathfrak{F}\varphi_{k}\!+\!\mathrm{i}\varphi_{k}\parallel$
1 $2.0\!\times\!10^{-10}$ 2 1.4142 1.4142
2 1.4142 1.4142 2 $1.2\!\times\!10^{-10}$
3 2 $6.3\!\times\!10^{-11}$ 1.4142 1.4142
4 1.4142 1.4142 $1.3\!\times\!10^{-12}$ 2
5 $2.7\!\times\!10^{-13}$ 2 1.4142 1.4142
6 1.4142 1.4142 2 $2.3\!\times\!10^{-14}$
7 2 $1.3\!\times\!10^{-14}$ 1.4142 1.4142
8 1.4142 1.4142 $3.4\!\times\!10^{-15}$ 2
9 $1.0\!\times\!10^{-15}$ 2 1.4142 1.4142
10 1.4142 1.4142 2 $8.3\!\times\!10^{-16}$
11 2 $8.4\!\times\!10^{-16}$ 1.4142 1.4142
If we compare the continuous and discrete version of the Fourier transform, then we arrive at the conclusion that, in the discrete case, the role of Planck constant is played by the dimension $d$ of the Hilbert space. Consequently, the discrete version of the relation (22) is
$\hat{\mathfrak{q}},\hat{\mathfrak{p}}]={\rm i}\mbox{\small$\frac{d}{2\pi}$}$ (25)
One can directly see that this relation is not satisfied, but by computing the eigenvalues of the operator
$\begin{array}[]{l}\hat{\mathfrak{C}}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},\quad\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\end{array}$ (26)
with the matrix elements
$\begin{array}[]{l}\langle n|\hat{\mathfrak{C}}|m\rangle\!=\!\frac{1}{d}\sum\limits_{k=-s}^{s}k(n\!-\!m)\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}k(n-m)]-{\rm i}\mbox{\small$\frac{d}{2\pi}$}\delta_{nm}\end{array}$ (27)
for different values of $d$, one arrives at the conclusion that the eigenvalues $\lambda$ with $|\lambda|<0.001$ represent $36\%$ in the case d=11d\!=\!11, represent $64\%$ in the case d=31d\!=\!31, represent $77\%$ in the case d=61d\!=\!61 and $84\%$ in the case d=101d\!=\!101 (see Table I). So, for $d$ large enough, a significant part of the eigenvalues of $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$ are almost null.
For any $\varepsilon\!>\!0$, there exists $d_{\varepsilon}$ such that the set
[…]
is non-empty for $d\!=\!\mathrm{dim}(\mathcal{H})\!>\!d_{\varepsilon}$. If $\varepsilon\!>\!0$ is small enough, then we can consider that
$\begin{array}[]{l}[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\frac{d}{2\pi}\psi\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon}.\end{array}$ (29)
If the states $\psi_{1}$, $\psi_{2}$, … , $\psi_{k}$ from $\mathcal{S}_{\varepsilon}$ form an orthonormal system and $\psi\!=\!\sum\limits_{m=1}^{k}\alpha_{m}\psi_{m}$ is such that $\psi||\!=\!1$, then
[…]
that is $\psi\!\in\!\mathcal{S}_{\varepsilon\sqrt{k}}$.
Table 3: Case d=11d\!=\!11: for certain $\kappa$, the Gaussian $\mathbf{g}_{\kappa}$ is mainly a linear combination of $\varphi_{1},$ $\varphi_{2},\,\varphi_{3},\,\varphi_{4},\,\varphi_{5}$.
$\kappa\!=\!1$ $\kappa\!=\!2$ $\kappa\!=\!1/2$ $\kappa\!=\!3$ $\kappa\!=\!1/3$
$|\langle\varphi_{0}|\mathbf{g}_{\kappa}\rangle|$ 0.999968 0.970268 0.970268 0.926811 0.926811
$|\langle\varphi_{1}|\mathbf{g}_{\kappa}\rangle|$ $8.79611\!\times\!10^{-11}$ $8.56453\!\times\!10^{-11}$ $8.50632\!\times\!10^{-11}$ $8.19559\!\times\!10^{-11}$ $8.11208\!\times\!10^{-8}$
$|\langle\varphi_{2}|\mathbf{g}_{\kappa}\rangle|$ $3.14918\!\times\!10^{-11}$ 0.228139 0.228139 0.327299 0.327299
$|\langle\varphi_{3}|\mathbf{g}_{\kappa}\rangle|$ $3.83482\!\times\!10^{-13}$ $3.8548\!\times\!10^{-13}$ $3.58495\!\times\!10^{-13}$ $74436\!\times\!10^{-13} 3.3$$36051\!\times\!10^{-13}$
$|\langle\varphi_{4}|\mathbf{g}_{\kappa}\rangle|$ $0.00798631$ 0.0736351 0.0736351 0.151392 0.151392
$|\langle\varphi_{5}|\mathbf{g}_{\kappa}\rangle|$ $8.61761\!\times\!10^{-15}$ $9.0834\!\times\!10^{-15}$ $8.25462\!\times\!10^{-15}$ $9.15919\!\times\!10^{-15}$ $8.20907\!\times\!10^{-15}$
$|\langle\varphi_{6}|\mathbf{g}_{\kappa}\rangle|$ $2.78121\!\times\!10^{-15}$ 0.0288995 0.0288995 0.0830314 0.0830314
$|\langle\varphi_{7}|\mathbf{g}_{\kappa}\rangle|$ $1.82021\!\times\!10^{-16}$ $1.7344\!\times\!10^{-16}$ $2.47453\!\times\!10^{-16}$ $1.32587\!\times\!10^{-16}$ $3.49103\!\times\!10^{-16}$
$|\langle\varphi_{8}|\mathbf{g}_{\kappa}\rangle|$ $0.000416018$ 0.0146752 0.0146752 0.0549712 0.0549712
$|\langle\varphi_{9}|\mathbf{g}_{\kappa}\rangle|$ $5.0058\!\times\!10^{-17}$ $4.89024\!\times\!10^{-17}$ $6.26338\!\times\!10^{-17}$ $4.68507\!\times\!10^{-17}$ $7.58437\!\times\!10^{-17}$
$|\langle\varphi_{10}|\mathbf{g}_{\kappa}\rangle|$ $1.98563\!\times\!10^{-16}$ 0.00777324 0.00777324 0.0325688 0.0325688
By using the relation
$\begin{array}[]{r}(\hat{\mathfrak{F}}\hat{\mathfrak{q}}\hat{\mathfrak{F}}^{\dagger})\psi(n)\!=\!\frac{1}{d}\sum\limits_{m,k=-s}^{s}\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}m(k-n)]\,m\,\psi(k)\\ =\!\frac{1}{d}\sum\limits_{m,k=-s}^{s}\mathrm{Exp}[-\frac{2\pi\mathrm{i}}{d}m(k-n)]\,(-m)\,\psi(k)\\ =\!-(\hat{\mathfrak{F}}^{\dagger}\hat{\mathfrak{q}}\hat{\mathfrak{F}})\psi(n)\!=\!-\hat{\mathfrak{p}}\psi(n)\end{array}$
we get $\hat{\mathfrak{F}}\hat{\mathfrak{q}}\hat{\mathfrak{F}}^{\dagger}\!=\!-\hat{\mathfrak{p}}$ and
[…]
that is
$\hat{\mathfrak{F}}[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\hat{\mathfrak{F}}$. (31)
From the relation
[…]
it follows that any set $\mathcal{S}_{\varepsilon}$ is Fourier invariant, that is
$\psi\!\in\!\mathcal{S}_{\varepsilon}\quad\Rightarrow\quad\hat{\mathfrak{F}}\psi\!\in\!\mathcal{S}_{\varepsilon}.$ (33)
Let $\lambda_{0}$, $\lambda_{1}$, …, $\lambda_{d-1}$ be the eigenvalues of $\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$, considered in the increasing order of their modulus ($|\lambda_{0}|\leq|\lambda_{1}|\leq...\leq|\lambda_{d-1}|$), and let $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ be the corresponding eigenfunctions, that is
$\hat{\mathfrak{C}}\varphi_{k}\!=\!\lambda_{k}\ \varphi_{\kappa}.$ (34)
The relation
$\hat{\mathfrak{C}}\varphi_{k}\!=\!\lambda_{k}\,\varphi_{\kappa}\quad\Rightarrow\quad\hat{\mathfrak{C}}\hat{\mathfrak{F}}\varphi_{k}\!=\!\hat{\mathfrak{F}}\hat{\mathfrak{C}}\varphi_{k}\!=\!\lambda_{k}\ \hat{\mathfrak{F}}\,\varphi_{\kappa}$
and the numerical data (see Table 2) suggest that the states $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ are also eigenstates of $\hat{\mathfrak{F}}$, namely
$\hat{\mathfrak{F}}\varphi_{n}\!=\!(-\mathrm{i})^{n}\ \varphi_{n}.$ (35)
Particularly, the discrete Fourier transform admits the spectral resolution
$\hat{\mathfrak{F}}\!=\!\sum_{n=0}^{d-1}(-\mathrm{i})^{n}\ |\varphi_{n}\rangle\langle\varphi_{n}|.$ (36)
For any $\varepsilon>0$ and $d$ high enough, some of the states $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{d-1}$ belong to $\mathcal{S}_{\varepsilon}$. For example (see Table 1), in the case d=31d\!=\!31, the states $\varphi_{0}$, $\varphi_{1}$, … , $\varphi_{19}$ belong to $\mathcal{S}_{0.001}$.
VI A discrete version of the uncertainty relation
A direct consequence of (29) is
$|\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|\approx\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon}.$ (37)
From the Robertson-Schrödinger uncertainty relation, it follows that the inequality
$\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\frac{|\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi|}{2}\quad\mbox{for}\ \ \psi\!\in\!\mathcal{S}_{\varepsilon},$ (38)
that is
$\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ (39)
is approximately satisfied in $\mathcal{S}_{\varepsilon}$.
VII Discrete-variable Gaussian functions
Figure 1: Discrete Gaussian functions: $\mathbf{g}_{3}$ and $\mathbf{g}_{1/3}$ in the case d=31d\!=\!31.
For any $\kappa\!\in\!(0,\infty)$, the Gaussian function $g_{\kappa}:\mathbb{R}\rightarrow\mathbb{R}$,
$\begin{array}[]{l}g_{\kappa}(q)\!=\!\mathrm{Exp}[-\frac{\kappa\pi}{h}q^{2}],\end{array}$ (40)
satisfies the relation
$\hat{F}[g_{\kappa}]\!=\!\mbox{\small$\frac{1}{\sqrt{\kappa}}$}\,g_{1/\kappa}.$ (41)
The corresponding discrete-variable Gaussian function
$\begin{array}[]{l}\mathfrak{g}_{\kappa}\!:\!\{-s,\!-s\!+\!\!1,...,s\!-\!\!1,\!s\}\!\rightarrow\!\mathbb{R},\\[2.84526pt] \mathfrak{g}_{\kappa}(n)=\sum\limits_{\alpha=-\infty}^{\infty}\mathrm{Exp}[-\frac{\kappa\pi}{d}(n+\alpha d)^{2}],\end{array}$ (42)
satisfies the relation [2]
$\hat{\mathfrak{F}}[\mathfrak{g}_{\kappa}]\!=\!\mbox{\small$\frac{1}{\sqrt{\kappa}}$}\ \mathfrak{g}_{1/\kappa}$ (43)
similar to (41). It is a particular case of Jacobi $\theta$ function [2, 7]. Because $\hat{\mathfrak{F}}$ is a unitary transform, the corresponding normalized discrete Gaussian functions (see Figure 1)
[…]
satisfy the relation
$\hat{\mathfrak{F}}[\mathbf{g}_{\kappa}]\!=\!\mathbf{g}_{1/\kappa}.$ (45)
For $d$ large enough and small $\delta>0$, there exists $\eta>0$ such that, for $\kappa\!\in\!(\frac{1}{\eta},\eta)$, the most significant coordinates of $\mathbf{g}_{\kappa}$ in the eigenbasis $\{\varphi_{1},\,\varphi_{2},\,...,\,\varphi_{d}\}$ of $\hat{\mathfrak{C}}\!=\![\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}$ are those corresponding to the functions $\varphi_{k}$ with $|\lambda_{k}|<\delta$ (see Table 3). We can consider that
$\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{for}\ \mathbf{g}_{\kappa}\ \mbox{with}\ \ \kappa\!\in\!(\mbox{\small$\frac{1}{\eta}$},\eta)$ (46)
(see Tables 3, 4 and Figure 2). From (35), we get
$|\langle\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\theta_{n}\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\mathfrak{F}\varphi_{n}|\mathbf{g}_{\kappa}\rangle|\!=\!|\langle\varphi_{n}|\mathbf{g}_{1/\kappa}\rangle|,$ (47)
in agreement with the numerical data from Table 3. Since $\hat{\mathfrak{F}}$ is a unitary transform, we get
[…]
for any $\psi\!\in\!\mathcal{H}$. Particularly, we have
$\parallel\mathfrak{C}\mathbf{g}_{\kappa}\parallel=\parallel\mathfrak{C}\mathbf{g}_{1/\kappa}\parallel,$ (49)
in agreement with the numerical data from Table 4.
By using (31) and (45), we get
$\langle\mathbf{g}_{1/\kappa}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{1/\kappa}\rangle\!=\!\langle\hat{\mathfrak{F}}\mathbf{g}_{\kappa}|\hat{\mathfrak{F}}[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{\kappa}\rangle\!=\!\langle\mathbf{g}_{\kappa}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{\kappa}\rangle,$ (50)
in agreement with the numerical data from Table 5.
Table 4: $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\mathbf{g}_{\kappa}\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\mathbf{g}_{\kappa}$ is satisfied for certain $\mathbf{g}_{\kappa}$.
$\kappa$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\mathbf{g}_{\kappa}||$
Case $d=11$ Case $d=31$
1 0.0022697 $2.68152\!\times\!10^{-9}$
2 0.283112 0.000670943
1/2 0.283112 0.000670943
3 0.0383181
1/3 0.0383181
Table 5: $|\langle\mathbf{g}_{\kappa}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{\kappa}\rangle|\approx\frac{d}{2\pi}$ is satisfied for certain $\mathbf{g}_{\kappa}$.
$\kappa$ $|\ |\langle\mathbf{g}_{\kappa}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{\kappa}\rangle|-\frac{d}{2\pi}\ |$
Case $d=11$ Case $d=31$
1 $1.21005\!\times\!10^{-6}$ $8.88178\!\times\!10^{-16}$
2 0.00350868 $2.47852\!\times\!10^{-9}$
1/2 0.00350868 $2.47852\!\times\!10^{-9}$
3 0.0552824 $7.39351\!\times\!10^{-6}$
1/3 0.0552824 $7.39351\!\times\!10^{-6}$
4 0.204735 0.000383435
1/4 0.204735 0.000383435
5 0.00397916
1/5 0.00397916
6 0.0185731
1/6 0.0185731
7 0.0550503
1/7 0.0550503
8 0.122973
1/8 0.122973
In conclusion, if the dimension $d$ of the Hilbert space describing the quantum system (qudit) is large enough, there exists a family $\mathcal{G}_{\delta}\!=\!\left\{\,\mathbf{g}_{\kappa}\ |\ \kappa\!\in\!(\mbox{\small$\frac{1}{\eta}$},\eta)\ \right\}$ of discrete-variable Gaussian states $\mathbf{g}_{\kappa}$ depending on a continuous parameter $\kappa\!\in\!(\mbox{\small$\frac{1}{\eta}$},\eta)$ such that the relations
$\hat{\mathfrak{q}},\hat{\mathfrak{p}}]={\rm i}\mbox{\small$\frac{d}{2\pi}$}\quad\mbox{and}\quad\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ (51)
are approximately satisfied for any $\mathbf{g}_{\kappa}\!\in\!\mathcal{G}_{\delta}$ (See Figure 2 and Figure 3).
Figure 2: Relation $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}$ in Gaussian case : $\parallel([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi})\mathbf{g}_{k}||$ for d=11d\!=\!11 and d=31d\!=\!31.
Figure 3: Relation $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ in Gaussian case : $|\ |\langle\mathbf{g}_{k}|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\mathbf{g}_{k}\rangle|-\frac{d}{2\pi}\ |$ for d=11d\!=\!11 and d=31d\!=\!31.
Any desired precision can be obtained by increasing the dimension $d$ of the Hilbert space and use of an adequate interval $(\mbox{\small$\frac{1}{\eta}$},\eta)$. In addition, the family $\mathcal{G}_{\delta}$ is Fourier invariant:
$\mathbf{g}_{\kappa}\!\in\!\mathcal{G}_{\delta}\quad\Rightarrow\quad\hat{\mathfrak{F}}[\mathbf{g}_{\kappa}]\!=\!\mathbf{g}_{1/\kappa}\!\in\!\mathcal{G}_{\delta}.$ (52)
VIII Discrete creation operator
The operators
[…]
can be regarded as a discrete version of the creation and annihilation operators
[…]
used in the continuous variable case. For small $\varepsilon\!>\!0$ and any $\psi\!\in\!\mathcal{S}_{\varepsilon}$, we have
$\hat{\mathfrak{a}},\hat{\mathfrak{a}}^{\dagger}]\!\approx\!1.$ (55)
It is known that the Hamiltonian of the harmonic oscillator can be written as
$\hat{H}\!=\!\hat{a}^{\dagger}\hat{a}\!+\!\frac{1}{2},$ (56)
and its eigenfunctions are
$\Psi_{n}\!=\!\frac{1}{\sqrt{n!}}(\hat{a}^{\dagger})^{n}\Psi_{0}$ (57)
where $\Psi_{0}(q)\!=\!\frac{1}{\sqrt[4]{\pi}}\mathrm{Exp}[-\frac{1}{2}q^{2}]$ is the vacuum state. By following the analogy with the continuous case, we define
[…]
where $\mathbf{g}\!\equiv\!\mathbf{g}_{1}$ is the discrete version of the vacuum state.
By direct computation, we can remark that, for small $\varepsilon\!>\!0$ and $d$ high enough, some of the states $\psi_{C,n}$ belong to $\mathcal{S}_{\varepsilon}$. For example (see Table 6), in the case d=31d\!=\!31, the states $\psi_{C,0}$, $\psi_{C,1}$, … , $\psi_{C,8}$ belong to $\mathcal{S}_{0.001}$.
Table 6: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain states $\psi_{C,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{C,0}$ $2.68152\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$
$\psi_{C,1}$ $3.67539\!\times\!10^{-8}$ 0
$\psi_{C,2}$ $4.93488\!\times\!10^{-7}$ $5.32907\!\times\!10^{-15}$
$\psi_{C,3}$ $2.24984\!\times\!10^{-6}$ $8.88178\!\times\!10^{-15}$
$\psi_{C,4}$ $9.70535\!\times\!10^{-6}$ $2.67342\!\times\!10^{-13}$
$\psi_{C,5}$ 0.000035908 $3.03935\!\times\!10^{-12}$
$\psi_{C,6}$ 0.000125868 $3.62173\!\times\!10^{-11}$
$\psi_{C,7}$ 0.000281649 $1.07889\!\times\!10^{-10}$
$\psi_{C,8}$ 0.000740305 $5.47327\!\times\!10^{-10}$
$\psi_{C,9}$ 0.00153664 $1.43238\!\times\!10^{-9}$
$\psi_{C,10}$ 0.00357292 $1.30984\!\times\!10^{-10}$
IX Discrete version of the harmonic oscillator
The operator $\hat{\mathfrak{H}}\!$$\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\hat{\mathfrak{H}}\!=\!\frac{\hat{\mathfrak{p}}^{2}\!+\!\hat{\mathfrak{q}}^{2}}{2}$ (59)
can be regarded as a discrete version of the Hamiltonian of the quantum harmonic oscillator
$\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}.$ (60)
For any small $\varepsilon\!>\!0$, some of the eigenstates of $\hat{\mathfrak{H}}$ belong to $\mathcal{S}_{\varepsilon}$. For example (see Table 7), in the case d=31d\!=\!31, the eigenstates $\psi_{O,0}$, $\psi_{O,1}$, … , $\psi_{O,7}$ belong to $\mathcal{S}_{0.001}$. They form an orthonormal system.
Table 7: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are satisfied for certain states $\psi_{O,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{O,0}$ $2.73402\!\times\!10^{-9}$ $5.32907\!\times\!10^{-15}$
$\psi_{O,1}$ $2.31647\!\times\!10^{-8}$ $1.77636\!\times\!10^{-15}$
$\psi_{O,2}$ $3.54883\!\times\!10^{-7}$ $4.44089\!\times\!10^{-15}$
$\psi_{O,3}$ $4.8082\!\times\!10^{-7}$ $1.15463\!\times\!10^{-14}$
$\psi_{O,4}$ $5.27306\!\times\!10^{-6}$ $4.70735\!\times\!10^{-13}$
$\psi_{O,5}$ $0.000018252$ $4.30322\!\times\!10^{-12}$
$\psi_{O,6}$ $0.000167857$ $1.50534\!\times\!10^{-10}$
$\psi_{O,7}$ $0.000127787$ $7.24515\!\times\!10^{-10}$
$\psi_{O,8}$ $0.0010158$ $1.59893\!\times\!10^{-8}$
$\psi_{O,9}$ $0.00233029$ $6.25305\!\times\!10^{-8}$
$\psi_{O,10}$ $0.0176262$ $1.48606\!\times\!10^{-6}$
X Discrete coherent states
The discrete coherent states [2, 3] are
$\begin{array}[]{l}|n,k\rangle\!=\!\frac{1}{\sqrt{d}}D(n,k)|\mathbf{g}\rangle,\end{array}$ (61)
where $n,k\!\in\!\{-\!s,\!-s\!+\!1,...,s\!-\!1,s\}$ and $D(n,k)\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H},$
$\begin{array}[]{l}D(n,k)\psi(m)=\mathrm{Exp}[-\frac{\pi\mathrm{i}}{d}nk]\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}km]\,\psi(m\!-\!n),\end{array}$ (62)
are the displacement operators. By using the relations
$\begin{array}[]{r}\parallel\!\mathfrak{C}|n,k\rangle\!\parallel^{2}=\!\!\!\sum\limits_{m=-s}^{s}\left|\frac{1}{d}\right.\!\!\sum\limits_{j=-s}^{s}\sum\limits_{\ell=-s}^{s}\!\!j(m\!-\!\ell)\,\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}j(m-\ell)]\\ \times\mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}k\ell]\,\mathbf{g}(\ell\!-\!n)\left.-\,\mathrm{i}\frac{d}{2\pi}\ \mathrm{Exp}[\frac{2\pi\mathrm{i}}{d}km]\ \mathbf{g}(m\!-\!n)\right|^{2},\end{array}$ (63)
$|\bar{z}|\!=\!|z|$, $\mathbf{g}(-n)\!=\!\mathbf{g}(n)$ and the changes of parameters $n\!\mapsto\!-n$, $k\!\mapsto\!-k$, $m\!\mapsto\!-m$, one can prove that
$\parallel\!\mathfrak{C}|n,k\rangle\!\!\parallel=\parallel\!\mathfrak{C}|-n,k\rangle\!\!\parallel=\parallel\!\mathfrak{C}|n,-k\rangle\!\!\parallel=\parallel\!\mathfrak{C}|-n,-k\rangle\!\!\parallel.$ (64)
For any small $\varepsilon\!>\!0$, some of the discrete coherent states $|j,k\rangle$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 8), in the case d=31d\!=\!31, the coherent states $|0,0\rangle$, $|0,\pm 1\rangle$, $|0,\pm 2\rangle$, $|0,\pm 3\rangle$, $|0,\pm 4\rangle$, $|\pm 1,0\rangle$, $|\pm 1,\pm 1\rangle$, $|\pm 1,\pm 2\rangle$, $|\pm 1,\pm 3\rangle$, $|\pm 1,\pm 4\rangle$, $|\pm 2,0\rangle$, $|\pm 2,\pm 1\rangle$, $|\pm 2,\pm 2\rangle$, $|\pm 2,\pm 3\rangle$, $|\pm 2,\pm 4\rangle$, $|\pm 3,0\rangle$, $|\pm 3,\pm 1\rangle$, $|\pm 3,\pm 2\rangle$, $|\pm 3,\pm 3\rangle$, $|\pm 3,\pm 4\rangle$, $|\pm 4,0\rangle$, $|\pm 4,\pm 1\rangle$, $|\pm 4,\pm 2\rangle$, $|\pm 4,\pm 3\rangle$, $|\pm 4,\pm 4\rangle$ belong to $\mathcal{S}_{0.001}$.
Table 8: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain states $|j,k\rangle$.
$|n,k\rangle$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)|n,k\rangle||$ $|\ |\langle n,k|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|n,k\rangle|-\frac{d}{2\pi}\ |$ $|n,k\rangle$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)|n,k\rangle||$ $|\ |\langle n,k|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|n,k\rangle|-\frac{d}{2\pi}\ |$
$|0,0\rangle$ $2.68152\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$ $|3,0\rangle$ $8.88946\!\times\!10^{-6}$ $2.70894\!\times\!10^{-13}$
$|0,1\rangle$ $3.53522\!\times\!10^{-8}$ $1.77636\!\times\!10^{-15}$ $|3,1\rangle$ $8.88748\!\times\!10^{-6}$ $2.70894\!\times\!10^{-13}$
$|0,2\rangle$ $6.06924\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$ $|3,2\rangle$ $9.21918\!\times\!10^{-6}$ $2.80664\!\times\!10^{-13}$
$|0,3\rangle$ $8.88946\!\times\!10^{-6}$ $2.73559\!\times\!10^{-13}$ $|3,3\rangle$ $0.000011576$ $5.21361\!\times\!10^{-13}$
$|0,4\rangle$ $0.000105793$ $4.01679\!\times\!10^{-11}$ $|3,4\rangle$ $0.000107044$ $4.00791\!\times\!10^{-11}$
$|0,5\rangle$ $0.00102558$ $3.9099\!\times\!10^{-9}$ $|3,5\rangle$ $0.00104948$ $3.91981\!\times\!10^{-9}$
$|1,0\rangle$ $3.53522\!\times\!10^{-8}$ $8.88178\!\times\!10^{-16}$ $|4,0\rangle$ $0.000105793$ $4.01696\!\times\!10^{-11}$
$|1,1\rangle$ $3.53347\!\times\!10^{-8}$ $2.66454\!\times\!10^{-15}$ $|4,1\rangle$ $0.000106023$ $4.01714\!\times\!10^{-11}$
$|1,2\rangle$ $6.2644\!\times\!10^{-7}$ 0 $|4,2\rangle$ $0.000106471$ $4.01306\!\times\!10^{-11}$
$|1,3\rangle$ $8.88748\!\times\!10^{-6}$ $2.74447\!\times\!10^{-13}$ $|4,3\rangle$ $0.000107044$ $4.00791\!\times\!10^{-11}$
$|1,4\rangle$ $0.000106023$ $4.01732\!\times\!10^{-11}$ $|4,4\rangle$ $0.000186876$ $9.42757\!\times\!10^{-11}$
$|1,5\rangle$ $0.00102781$ $3.90986\!\times\!10^{-9}$ $|4,5\rangle$ $0.00100292$ $3.78383\!\times\!10^{-9}$
$|2,0\rangle$ $6.06924\!\times\!10^{-7}$ $1.77636\!\times\!10^{-15}$ $|5,0\rangle$ $0.00102558$ $3.9099\!\times\!10^{-9}$
$|2,1\rangle$ $6.2644\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$ $|5,1\rangle$ $0.00102781$ $3.90986\!\times\!10^{-9}$
$|2,2\rangle$ $6.28767\!\times\!10^{-7}$ $2.66454\!\times\!10^{-15}$ $|5,2\rangle$ $0.00103463$ $3.91002\!\times\!10^{-9}$
$|2,3\rangle$ $9.21918\!\times\!10^{-6}$ $2.77112\!\times\!10^{-13}$ $|5,3\rangle$ $0.00104948$ $3.91981\!\times\!10^{-9}$
$|2,4\rangle$ $0.000106471$ $4.01323\!\times\!10^{-11}$ $|5,4\rangle$ $0.00100292$ $3.78383\!\times\!10^{-9}$
$|2,5\rangle$ $0.00103463$ $3.91002\!\times\!10^{-9}$ $|5,5\rangle$ $0.00170788$ $8.59234\!\times\!10^{-9}$
XI Discrete coherent state quantization
In the continuous case the standard coherent states
$|q,p\rangle\!=\!D(q,p)|g\rangle,$ (65)
where
$\begin{array}[]{l}g(q)\!=\!\frac{1}{\sqrt[4]{\pi}}\,\mathrm{Exp}[-\frac{1}{2}q^{2}]\end{array}$ (66)
is the vacuum state, satisfy the resolution of the identity
$\begin{array}[]{l}\frac{1}{2\pi}\int\!\!\!\int\limits_{\mathbb{R}^{2}}|q,p\rangle\langle q,p|\,dqdp\!=\!\mathbb{I}_{L^{2}(\mathbb{R})}.\end{array}$ (67)
To any function $f\!:\!\mathbb{R}^{2}\!\rightarrow\mathbb{C}$ defined on the phase space $\mathbb{R}^{2}$, and chosen such that the integral is convergent, one can associate the operator
$\begin{array}[]{l}A_{f}\!=\!\frac{1}{2\pi}\int\!\!\!\int\limits_{\mathbb{R}^{2}}f(q,p)\,|q,p\rangle\langle q,p|\,dqdp.\end{array}$ (68)
The map $f\!\mapsto A_{f}$ is called coherent state quantization [3, 15, 16] . One can prove that [15]
$\begin{array}[]{lll}A_{f}\!=\!\hat{q}&\mbox{for}&f(q,p)=q,\\ A_{f}\!=\!\hat{p}&\mbox{for}&f(q,p)=p,\\ A_{f}\!=\!\frac{1}{2}(\hat{p}^{2}\!+\!\hat{q}^{2})\!+\!\frac{1}{2}&\mbox{for}&f(q,p)=\frac{1}{2}({p}^{2}\!+\!{q}^{2}).\end{array}$ (69)
Because the discrete coherent states satisfy
$\sum\limits_{n,k=-s}^{s}|n,k\rangle\langle n,k|\!=\!\mathbb{I},$ (70)
by following the analogy with the continuous case, one can associate the operator $\mathfrak{A}_{f}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$\begin{array}[]{l}\mathfrak{A}_{f}\!=\!\sum\limits_{n,k=-s}^{s}f(n,k)\,|n,k\rangle\langle n,k|\end{array}$ (71)
to each function
$f\!:\!\{-s,-s\!+\!1,...,s\}\!\times\!\{-s,-s\!+\!1,...,s\}\!\rightarrow\mathbb{C}$ (72)
defined on the discrete phase space. By using the discrete coherent state quantization $f\!\mapsto\!\mathfrak{A}_{f}$, we obtain for $f(n,k)\!=\!\frac{1}{2}(n^{2}\!+\!k^{2})$ the operator
$\begin{array}[]{l}\mathfrak{H}_{q}\!=\!-\frac{1}{2}\!+\!\sum\limits_{n,k=-s}^{s}\frac{n^{2}\!+\!k^{2}}{2}\,|n,k\rangle\langle n,k|\end{array}$ (73)
which can be regarded as a discrete version of the harmonic oscillator Hamiltonian. For any small $\varepsilon\!>\!0$, some of the eigenfunctions $\psi_{Q,n}$ of $\mathfrak{H}_{q}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 9), in the case d=31d\!=\!31, the states $\psi_{Q,0}$, $\psi_{Q,1}$, … , $\psi_{Q,7}$, where the states have been considered in the increasing order of eigenvalues, belong to $\mathcal{S}_{0.001}$.
Table 9: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain states $\psi_{Q,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{Q,0}$ $3.95539\!\times\!10^{-9}$ $1.77636\!\times\!10^{-15}$
$\psi_{Q,1}$ $3.45774\!\times\!10^{-8}$ $2.66454\!\times\!10^{-15}$
$\psi_{Q,2}$ $5.48854\!\times\!10^{-7}$ $4.44089\!\times\!10^{-15}$
$\psi_{Q,3}$ $6.77894\!\times\!10^{-7}$ $2.4869\!\times\!10^{-14}$
$\psi_{Q,4}$ $7.94567\!\times\!10^{-9}$ $9.82325\!\times\!10^{-13}$
$\psi_{Q,5}$ $0.0000278873$ $8.32134\!\times\!10^{-12}$
$\psi_{Q,6}$ $0.00027139$ $3.52218\!\times\!10^{-10}$
$\psi_{Q,7}$ $0.000187143$ $1.42712\!\times\!10^{-9}$
$\psi_{Q,8}$ $0.00162965$ $3.83563\!\times\!10^{-8}$
$\psi_{Q,9}$ $0.00376062$ $1.33373\!\times\!10^{-7}$
$\psi_{Q,10}$ $0.0310428$ $4.15222\!\times\!10^{-6}$
XII Discrete annihilation operator
It is known that, in the continuous-variable case, the standard coherent states $|q,p\rangle$ can be defined as eigenstates of the annihilation operator
$\begin{array}[]{l}\hat{a}\,|q,p\rangle\!=\!\sqrt{\frac{\pi}{h}}(q\!+\!\mathrm{i}p)|q,p\rangle.\end{array}$ (74)
By following the analogy, we consider the eigenstates $\psi_{A,n}$ of the discrete version $\hat{\mathfrak{a}}$ of $\hat{a}$. For any small $\varepsilon\!>\!0$, some of the eigenfunctions $\psi_{A,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 10), in the case d=31d\!=\!31, the eigenstates $\psi_{A,0}$, $\psi_{A,1}$, … , $\psi_{A,14}$, where the states have been considered in the increasing order of the number of sign oscillations, belong to $\mathcal{S}_{0.001}$.
h $\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$ $\psi_{A,0}$ $2.08125\!\times\!10^{-9}$ 0   $\psi_{A,1}$ $2.08125\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$ $\psi_{A,2}$ $2.08126\!\times\!10^{-9}$ $1.77636\!\times\!10^{-15}$ $\psi_{A,3}$ $1.41314\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$ $\psi_{A,4}$ $1.41314\!\times\!10^{-7}$ $8.88178\!\times\!10^{-16}$ $\psi_{A,5}$ $1.41314\!\times\!10^{-7}$ $1.77636\!\times\!10^{-15}$ $\psi_{A,6}$ $1.41314\!\times\!10^{-7}$ 0   $\psi_{A,7}$ $3.2588\!\times\!10^{-6}$ $2.57572\!\times\!10^{-14}$ $\psi_{A,8}$ $3.2588\!\times\!10^{-6}$ $2.66454\!\times\!10^{-14}$ $\psi_{A,9}$ $3.2588\!\times\!10^{-6}$ $2.57572\!\times\!10^{-14}$ $\psi_{A,10}$ $3.2588\!\times\!10^{-6}$ $2.75335\!\times\!10^{-14}$ $\psi_{A,11}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,12}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,13}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,14}$ 0.000746798 $1.2461\!\times\!10^{-11}$ $\psi_{A,15}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,16}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,17}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,18}$ 0.00169824 $5.93746\!\times\!10^{-9}$ $\psi_{A,19}$ 0.0381896 $2.72907\!\times\!10^{-6}$
Table 10: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied by certain states $\psi_{A,k}$.
XIII Mehta states
It is known that, in the continuous-variable case, the eigenstates of the quantum harmonic oscillator $\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}$ can be defined in terms of Hermite polynomials as $\Psi_{n}\!:\!(-\infty,\infty)\!\rightarrow\mathbb{R}$,
$\begin{array}[]{l}\Psi_{n}(q)\!=\!\frac{1}{\sqrt{n!\,2^{n}\sqrt{\pi}}}\,H_{n}(q)\ \mathrm{Exp}[-\frac{1}{2}q^{2}]\end{array}$ (75)
and satisfy the relation
$\begin{array}[]{l}\hat{H}\Psi_{n}\!=\!(n\!+\!\frac{1}{2})\Psi_{n}.\end{array}$ (76)
It is known that the Hermite-Gauss functions $\Psi_{n}$ are also eigenfunctions of the Fourier transform, namely
$\hat{F}\Psi_{n}\!=\!(-\mathrm{i})^{n}\Psi_{n}.$ (77)
In his article [7], Mehta has defined the discrete version $\psi_{M,n}\!:\!\{-s,-s\!+\!1,...,s\}\!\rightarrow\!\mathbb{R}$,
$\begin{array}[]{l}\psi_{M,n}(m)\!=\!\sum\limits_{\alpha=-\infty}^{\infty}H_{n}(m\!+\!\alpha d)\ \mathrm{Exp}[-\frac{1}{2}(m\!+\!\alpha d)^{2}]\end{array}$ (78)
of $\Psi_{n}$, and proved that it is an eigenfunction of the discrete Fourier transform,
$\hat{\mathfrak{F}}\psi_{M,n}\!=\!(-\mathrm{i})^{n}\psi_{M,n}.$ (79)
For any small $\varepsilon\!>\!0$, some states $\psi_{M,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 11), in the case d=31d\!=\!31, the functions $\psi_{M,0}$, $\psi_{M,1}$, … , $\psi_{M,9}$, after normalization, belong to $\mathcal{S}_{0.001}$.
Table 11: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain Mehta states $\psi_{M,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{M,0}$ $2.68152\!\times\!10^{-9}$ $8.88178\!\times\!10^{-16}$
$\psi_{M,1}$ $2.41961\!\times\!10^{-8}$ 0
$\psi_{M,2}$ $2.98432\!\times\!10^{-7}$ $2.66454\!\times\!10^{-15}$
$\psi_{M,3}$ $349482\!\times\!10^{-7}$ $1.1543\!\times\!10^{-14}$
$\psi_{M,4}$ $2.95378\!\times\!10^{-6}$ $1.55431\!\times\!10^{-13}$
$\psi_{M,5}$ $8.69175\!\times\!10^{-6}$ $9.28146\!\times\!10^{-13}$
$\psi_{M,6}$ 0.00005669 $1.78559\!\times\!10^{-11}$
$\psi_{M,7}$ 0.00003822 $5.83524\!\times\!10^{-11}$
$\psi_{M,8}$ 0.0002320 $7.42824\!\times\!10^{-10}$
$\psi_{M,9}$ 0.0004216 $2.07663\!\times\!10^{-9}$
$\psi_{M,10}$ 0.0022077 $2.6385\!\times\!10^{-8}$
XIV Harper states
In continuous-variable case, $\hat{p}\!=\!\hat{F}^{\dagger}\hat{q}\hat{F}$, and the harmonic oscillator Hamiltonian can be written as
$\hat{H}\!=\!\frac{\hat{p}^{2}\!+\!\hat{q}^{2}}{2}\!=\!\frac{\hat{p}^{2}\!+\!\hat{F}\hat{p}^{2}\hat{F}^{\dagger}}{2}.$ (80)
A discrete version of the quantum harmonic oscillator can be obtained by using for the differential operator $\frac{\mathrm{d}^{2}}{\mathrm{d}q^{2}}$ the finite-difference version $D^{2}\!:\!\mathcal{H}\!\rightarrow\!\mathcal{H}$,
$D^{2}\psi(n)\!=\!\psi(n\!-\!1)\!-\!2\psi(n)\!+\!\psi(n\!+\!1)$ (81)
and the relation
$\begin{array}[]{l}\hat{\mathfrak{F}}D^{2}\hat{\mathfrak{F}}^{\dagger}\psi(n)\!=\!\left(2\cos\frac{2\pi n}{d}\!-\!1\right)\psi(n).\end{array}$ (82)
The eigenfunctions $\psi_{H,n}\!:\!\{-s,-s\!+\!1,...,s,s\!+\!1\}\!\rightarrow\!\mathbb{R}$ of
$\hat{\mathfrak{H}}_{d}\!=\!-\frac{1}{2}(D^{2}\!+\!\hat{\mathfrak{F}}D^{2}\hat{\mathfrak{F}}^{\dagger})$ (83)
correspond to the normalized eigenvectors of the matrix (addition is modulo $d$)
$\begin{array}[]{l}\left(2\cos\frac{2\pi n}{d}\delta_{nk}\!-\!4\!+\!\delta_{k,n+1}\!+\!\delta_{k,n-1}\right)_{\!\!-s\leq n,k\leq s}\end{array}$ (84)
and are called Harper functions [8].
For any small $\varepsilon\!>\!0$, some states $\psi_{H,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 12), in the case d=31d\!=\!31, the eigenstates $\psi_{H,0}$, $\psi_{H,1}$, $\psi_{H,2}$ , $\psi_{H,3}$ belong to $\mathcal{S}_{0.001}$.
Table 12: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain Harper states $\psi_{H,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{H,0}$ $1.75165\!\times\!10^{-6}$ $4.08562\!\times\!10^{-14}$
$\psi_{H,1}$ 0.0000101661 $8.56204\!\times\!10^{-13}$
$\psi_{H,2}$ 0.000218006 $1.87562\!\times\!10^{-10}$
$\psi_{H,3}$ 0.000136764 $6.73229\!\times\!10^{-10}$
$\psi_{H,4}$ 0.00211415 $5.87102\!\times\!10^{-8}$
$\psi_{H,5}$ 0.00445289 $1.51256\!\times\!10^{-7}$
$\psi_{H,6}$ 0.0632087 $0.0000150119$
$\psi_{H,7}$ 0.0187208 $0.0000122191$
XV Kravchuk states
The Kravchuk polynomials $K_{m}$, defined by
$(1\!-\!X)^{s+k}(1\!+\!X)^{s-k}\!=\!\sum\limits_{m=-s}^{s}K_{m}(k)\,X^{s+m}$ (85)
are
$K_{m}(k)\!=\!\sum\limits_{n=0}^{s+m}(-1)^{n}\ C_{s+k}^{m}\ C_{s-k}^{s+m-n}$ (86)
and satisfy the relation
$\begin{array}[]{l}\frac{1}{2^{2s}}\sum\limits_{k=-s}^{s}C_{2s}^{s+k}\ K_{m}(k)\ K_{n}(k)\!=\!C_{2s}^{s+m}\ \delta_{mn},\end{array}$ (87)
where
$C_{m}^{n}\!=\!\left\{\begin{array}[]{cll}\frac{m!}{n!\ (m\!-\!n)!}&\mbox{for}&n\!\in\!\{0,1,...,m\}\\ 0&\mbox{for}&n\!\not\in\!\{0,1,...,m\}.\end{array}\right$. (88)
The Kravchuk functions [9, 10]
$\psi_{K,-s},\ \psi_{K,-s+1},\,...\,\ \psi_{K,s}:\{-s,-s\!+\!1,...,s\!-\!1,s\}\!\rightarrow\!\mathbb{R},$
$\psi_{K,m}(n)\!=\!\frac{1}{2^{s}}\sqrt{\frac{C_{2s}^{s+n}}{C_{2s}^{s+m}}}\ K_{m}(n),$ (89)
form an orthonormal basis in $\mathcal{H}$.
For any small $\varepsilon\!>\!0$, some states $\psi_{K,n}$ belong to $\mathcal{S}_{\varepsilon}$ if the Hilbert space of the quantum system has a dimension high enough. For example (see Table 13), in the case d=31d\!=\!31, the eigenstates $\psi_{K,-15}$, $\psi_{K,-14}$, $\psi_{K,-13}$, belong to $\mathcal{S}_{0.1}$.
Table 13: In the case d=31d\!=\!31, the relations $\hat{\mathfrak{q}},\hat{\mathfrak{p}}]\psi\approx{\rm i}\mbox{\small$\frac{d}{2\pi}$}\psi$ and $\Delta\hat{\mathfrak{q}}\ \Delta\hat{\mathfrak{p}}\geq\mbox{\small$\frac{d}{4\pi}$}$ are approximately satisfied for certain Kravchuk states $\psi_{K,k}$.
$\psi$ $\left([\hat{\mathfrak{q}},\hat{\mathfrak{p}}]-{\rm i}\frac{d}{2\pi}\right)\psi||$ $|\ |\langle\psi|[\hat{\mathfrak{q}},\hat{\mathfrak{p}}]|\psi\rangle|-\frac{d}{2\pi}\ |$
$\psi_{K,-15}$ 0.00188 $3.3511\!\times\!10^{-8}$
$\psi_{K,-14}$ 0.010478 $1.06622\!\times\!10^{-6}$
$\psi_{K,-13}$ 0.0435235 $0.0000165$
XVI Concluding remarks
In the case of a continuous-variable quantum system, the canonical position-momentum commutation relation is exactly satisfied in a subspace of the Hilbert space, and plays a fundamental role. The discrete version of this important relation is also satisfied in a subspace, but not exactly, and this happens only if the dimension of the Hilbert space is high enough. We have investigated the behaviour with respect to the discrete version of some important states used in the mathematical description of qudits, including discrete Gaussian states, discrete coherent states and discrete versions of the Hermite-Gauss states. We think that their property discovered in this way is important enough, and may have certain applications.
Quantum Information uses a continuous-variable description, but it is mainly limited to a very particular class of states (including the Gaussian states) and a very particular class of unitary transforms (including the Gaussian unitaries). The discrete-variable version of the commutation relation of the discrete-variable version of position and momentum operator is acceptably satisfied by certain particular states. We think that, some useful models involving only such particular states, can be built in the discrete-variable case by following the analogy with the continuous-variable case. A first attempt in this direction can be found in [4].